# PinRec: Outcome-Conditioned, Multi-Token Generative Retrieval for Industry-Scale Recommendation Systems

### v2

@@ A.1.1. HSTU Architecture
Though PinRec relies primarily on the well-known transformer architecture, we note that our technique is completely architecture-agnostic. Specifically, outcome conditioning and multi-token prediction can be applied regardless of the specific architecture (e.g., transformer blocks) or algorithms used (e.g., attention), allowing them to be generic for any future use case of generative retrieval despite architectural or algorithmic improvements. To demonstrate this, we reproduce significant gains over baselines using the HSTU architecture (Zhai et al., 2024), as opposed to the multi-head attention-based causal transformer. Note that we do not leverage the ranking task or any sequence sampling explored in (Zhai et al., 2024) as it is not relevant for our specific use case. For simplicity, we reproduce results on the outcome conditioned variant of PinRec only, as we believe similar results can be extrapolated to other variants.
@@ A.1.2. Ablating Negative Sampling
Following from Pancha et al. (2022), we perform ablations on negative sampling, as it is a critical component of the chosen sampled softmax loss function. Though we ultimately end up with similar conclusions as Pancha et al. (2022), we show these findings for maximum clarity regarding our design choices.
@@ A.1.4. Tail User and Item Analysis
In terms of new users, PinRec-UC/OC lift grid clicks and outbound clickthroughers (outbound CTs) in 9(a). Critically, the lift for outcome conditioning-based modeling is either neutral or positive with respect to both control and the unconditioned variant of PinRec, demonstrating that outcome conditioning does not suffer on such cohorts of users. In fact, across new users, OC performs better on outbound CTs.
@@ A.2. Visualization of Multi-Token Impact
In this section, we present some additional qualitative examples through visualizations of the retrieved examples for some sample user histories across the windowed multi-token and non-multi-token variants. Through these visualizations, we are able to demonstrate a reasonable and consistent qualitative difference in the types of pins retrieved, justifying the improved session-wide and sitewide online metrics for PinRec-{MT, OC}.

### v4

@@ A.1.1. HSTU Architecture
Though PinRec relies primarily on the well-known transformer architecture, we note that our technique is completely architecture-agnostic. Specifically, outcome conditioning and multi-token prediction can be applied regardless of the specific architecture (e.g., transformer blocks) or algorithms used (e.g., attention), allowing them to be generic for any future use case of generative retrieval despite architectural or algorithmic improvements. To demonstrate this, we reproduce significant gains over baselines using the HSTU architecture (HSTU24), as opposed to the multi-head attention-based causal transformer. Note that we do not leverage the ranking task or any sequence sampling explored in (HSTU24) as it is not relevant for our specific use case. For simplicity, we reproduce results on the outcome conditioned variant of PinRec only, as we believe similar results can be extrapolated to other variants.
@@ A.1.2. Ablating Negative Sampling
Following from Pinnerformer22, we perform ablations on negative sampling, as it is a critical component of the chosen sampled softmax loss function. Though we ultimately end up with similar conclusions as Pinnerformer22, we show these findings for maximum clarity regarding our design choices.
@@ A.1.4. Tail User and Item Analysis
In terms of new users, PinRec-UC/OC lift grid clicks and outbound clickthroughers (outbound CTs) in 10(a). Critically, the lift for outcome conditioning-based modeling is either neutral or positive with respect to both control and the unconditioned variant of PinRec, demonstrating that outcome conditioning does not suffer on such cohorts of users. In fact, across new users, OC performs better on outbound CTs.
@@
(a) PinRec-OC
(b) PinRec-{MT, OC}
Figure 8. Visualization of Pareto fronts for online lift % in Homefeed grid clicks versus repins for PinRec variants.
@@ A.2. Visualization of Multi-Token Impact
In this section, we present additional qualitative examples through visualizations of retrieved examples for sample user histories, comparing multi-token and non-multi-token variants. Through these visualizations, we are able to demonstrate a reasonable and consistent qualitative difference in the types of pins retrieved, justifying the improved session-wide and sitewide online metrics for PinRec-{MT, OC}.